OpenStudy (anonymous):

please Help!!!! 4.Determine two different values of “b” in x2 + bx + 30 so that the expression can be factored into the product of two binomials. Explain how you determined those values and show each factorization. Explain how your process would change if the expression was 2x2 + bx + 30.

OpenStudy (anonymous):

you know it must be in the form (x+c)(x+d), so you expand this to x^2+(c+d)x+cd, equate the coefficients with the equation you had originally so c+d=b and cd=30 by choosing an arbitrary value for c you can work out d, like so c=2, 2*d=30, so d=15, now you can work out b, 2+15=b so b=17. you can do this for as many values for c or d as you want. for 2x^2+bx+30 I would do the same except I would expand (2x+c)(x+d) instead, you would then do the same as before, equating coefficients and choosing arbitrary values for c or d.

OpenStudy (anonymous):

i found that on google hah, hope it helped some